OpenStudy (anonymous):

The equation of the line through point (- 4, 0) and parallel to the y-axis is

OpenStudy (anonymous):

answer has to be x = -4 . Or numerically assume (x,y) is a point on our line .If resulting line is || to Y-Axis then we know the slope m = \[\infty\] . Now, \[m = (y-0)\div(x+4)\]\[\frac{ 1 }{ m } = \frac{ x+4 }{ y-0 }\] \[\frac{ 1 }{ \infty } = \frac{ x+4 }{ y-0 }\]\[x + 4 = 0\]\[x = -4\]